SkinnyMinnie20:

Find the lateral area of the pyramid to the nearest whole number. 77 m2 85 m2 51 m2 47 m2

SkinnyMinnie20:

1 attachment
SkinnyMinnie20:

@Aeon

SkinnyMinnie20:

@cookiemonster

Vocaloid:

we have to do a little bit of math to figure out the height of the sides

Vocaloid:



Vocaloid:



Vocaloid:

according to pythagoreas, a^2 + b^2 = c^2 if 2^2 + 6^2 = c^2 then c = sqrt(40) = about 6.32

Vocaloid:

the lateral area is the sum of the four triangle faces so: 4(1/2)(6.32) = your answer